Please help with this probability question!

Started by katieTX
This forum made possible through the generous support of SDN members, donors, and sponsors. Thank you.
Get help with your application

Use all the free resources available to you from SDN: articles, guides, expert advising, forums discussions, and school research.

katieTX

Member
7+ Year Member
15+ Year Member
20+ Year Member
Advertisement - Members don't see this ad
I got this question wrong in Barron Model A Exam but they have not explaination for it. Please help.

A bowl contains 7 blue and 3 red marbles. What is the probability that 2 marbles selected at random from this bowl without replacement are both red?

I'm here waiting for reply 🙂 Mucho gracias1
 
Probability =
3 red out of 10 = 3/10 for first draw
2 red out of 9 = 2/9 for second draw

Total Probablity = 3/10 * 2/9 = 6/90 = 1/15